The length of a rectangle is 1 foot longer than its width. If the perimeter of the rectangle is 26 feet, find its length.
I've tried:
P=L+W+L+W=26
2L+2W=26

Thank you
Shanna

Answer by mananth(16946)

let the width be W feet
length is 1 ft more than the width= W+1 feet
..
Perimeter = 2(L+W)
2(W+W+1) = 26
2(2W+1)=26
/2
2W+1= 13
-1
2W=13-1
2W=12
/2
W=6 feet
Length = W+1 = 6+1 =7 feet.
...
C H E C K
6+6+7+7 = 26
...
m.ananth@hotmail.ca